# BPE: Behavioral Profiling Ensemble

### v1

@@ Abstract
Ensemble learning is widely recognized as a pivotal strategy for pushing the boundaries of predictive performance. Traditional static ensemble methods, such as Stacking, typically assign weights by treating each base learner as a holistic entity, thereby overlooking the fact that individual models exhibit varying degrees of competence across different regions of the instance space. To address this limitation, Dynamic Ensemble Selection (DES) was introduced. However, both static and dynamic approaches predominantly rely on the divergence among different models as the basis for integration. This inter-model perspective neglects the intrinsic characteristics of the models themselves and necessitates a heavy reliance on validation sets for competence estimation. In this paper, we propose the Behavioral Profiling Ensemble (BPE) framework, which introduces a novel paradigm shift. Unlike traditional methods, BPE constructs a “behavioral profile” intrinsic to each model and derives integration weights based on the deviation between the model’s response to a specific test instance and its established behavioral profile. Extensive experiments on both synthetic and real-world datasets demonstrate that the algorithm derived from the BPE framework achieves significant improvements over state-of-the-art ensemble baselines. These gains are evident not only in predictive accuracy but also in computational efficiency and storage resource utilization across various scenarios.
@@
journal: Nuclear Physics B
[…]
[inst2]organization=Yunnan University, city=Kunming, state=Yunnan, country=China
@@ 1 Introduction
In the field of machine learning, ensemble learning, which enhances predictive performance by combining multiple base learners, has become the dominant approach for addressing complex classification problems. Whether in data science competitions or high-stakes industrial applications such as medical diagnosis and financial risk control, ensemble models typically exhibit superior generalization capability and robustness compared to individual models by exploiting diversity among learners [1, 2]. Current mainstream ensemble strategies are generally categorized into two paradigms: global static ensembles, such as simple averaging, median averaging, or rank averaging [3, 4]; and meta-learning-based methods, such as Stacking, which learn the combination weights of base models by training a meta-classifier [5].
@@
Despite their substantial success, these methods predominantly rely on the assumption that a specific model possesses consistent superiority across the entire data distribution, thereby treating models as globally uniform. However, in practical scenarios, different models often manifest distinct strengths in various sample subspaces. This global perspective tends to overlook the local competence variances of the models. To address this issue, Dynamic Ensemble Selection (DES) has emerged [6]. Conventional DES methods typically follow a neighborhood-based validation paradigm: during the inference phase, they retrieve similar samples from a validation set and dynamically assign weights based on the models’ historical performance on these local neighbors [7, 8]. Nevertheless, such reliance on an external validation set poses challenges in deployment and limits generalization in certain contexts.
@@
To overcome these limitations, we propose a fundamental shift in perspective: transitioning from external validation matching to the exploitation of models’ intrinsic properties. This paradigm shift can be understood through an intuitive analogy: traditional DES algorithms resemble ”resume screening”, when a new task (test sample) arrives, the system reviews historical records to find individuals who performed well on similar past tasks (validation neighbors). The drawback is that for novel or outlier instances, the historical resume loses its reference value. In contrast, our proposed method functions as ”behavioral profiling” rather than focusing on whether the individual has encountered similar tasks before, we observe their inherent behavioral traits through instantaneous stress testing to designate the most reliable candidate.
@@
Driven by this motivation, this paper proposes the Behavioral Profiling Ensemble (BPE), a validation-free dynamic ensemble framework based on model ”personality” profiling. Specifically, we discard the traditional neighborhood search and instead introduce stochastic perturbations in the feature space to construct an instantaneous ”testing field” for each model. By observing the variations in the output distributions of different base models under perturbation (e.g., entropy increase), we can reveal the models’ inherent behavioral characteristics [9, 10]. This profiling effectively differentiates between models that are robustly confident and those that exhibit high uncertainty on the current sample, thereby serving as the basis for dynamic weighting. Extensive experiments on both simulation and real-world datasets demonstrate that BPE consistently outperforms traditional methods across numerous scenarios.
@@ 2.1 Ensemble Learning Theory and Uncertainty Measures
The core philosophy of ensemble learning lies in compensating for the inductive bias of a single learner by combining multiple hypotheses. The “No Free Lunch” theorem proposed by Wolpert and Macready [11] theoretically elucidates the necessity of multiple classifier systems, asserting that no single algorithm can perform optimally across all possible data distributions. To quantify the effectiveness of ensembles, Krogh and Vedelsby proposed the renowned “Error-Ambiguity Decomposition” theory [1], proving that the generalization error of an ensemble is strictly equal to the average error of the base models minus an ambiguity term representing model diversity. This theory establishes the dual objectives of ensemble learning: maximizing diversity among models while pursuing individual accuracy. From a statistical perspective, Dietterich further noted that ensemble methods effectively smooth decision boundaries by averaging multiple hypotheses, thereby significantly mitigating the dilemma of the bias-variance tradeoff for individual models [2].
To accurately allocate weights in dynamic ensembles, quantifying a single model’s “Predictive Confidence” on specific samples is crucial. Information entropy, as proposed by Shannon [12], serves as the cornerstone for measuring the uncertainty of probability distributions and is widely utilized to quantify classifier prediction confidence. Grandvalet and Bengio further demonstrated in semi-supervised learning that minimizing the entropy of the predictive distribution can effectively enhance model confidence on unlabeled data, establishing the theoretical connection between low entropy and high confidence [13]. In modern machine learning research, perturbation-based uncertainty estimation has become mainstream. Deep Ensembles, proposed by Lakshminarayanan et al. [10], demonstrated that epistemic uncertainty can be accurately estimated by ensembling multiple models. Similarly, Gal and Ghahramani demonstrated that enabling Dropout during the inference phase is equivalent to Bayesian approximation; by calculating predictive entropy through multiple stochastic forward passes, one can effectively capture the model’s intrinsic uncertainty [9]. Furthermore, Schapire’s margin theory indicates that maximizing the classification margin—manifested as significant differences in predictive probabilities—is key to enhancing generalization capability [14]. These studies provide a solid theoretical foundation for the proposed BPE framework, which constructs a behavioral profile of model entropy via training set perturbations.
@@ 2.2 Static Ensembles
Early ensemble strategies predominantly adopted a static perspective, where the performance weights of base models remain fixed across the entire feature space. Depending on the generation mechanism of the base learners, these strategies are primarily categorized into homogeneous and heterogeneous ensembles.
In homogeneous ensembles, the Bagging algorithm proposed by Breiman [15] constructs diversity via bootstrap sampling; its prominent representative, Random Forest [16], primarily enhances performance by mitigating variance. Conversely, Boosting paradigms aim to reduce bias by iteratively correcting residuals. Notable examples include the classic AdaBoost [17] and Gradient Boosting Machines (GBM) [18], along with highly optimized implementations such as XGBoost [19], LightGBM [20], and CatBoost [21].
In the realm of heterogeneous ensembles comprising diverse algorithms (e.g., SVC [22], KNN [23], and Logistic Regression [24]), the prevailing strategy involves static fusion rules. Kittler et al. systematized fundamental combination rules, including majority voting and simple averaging [3]. To address scale discrepancies in probability outputs, rank-based methods, such as the Borda Count applied by Ho et al. [4], mitigate these differences through rank aggregation. Furthermore, the Stacking framework [5] employs a meta-learner to learn the globally optimal combination of base models. Despite their widespread success, these static methods inherently overlook the locality of data distributions; a model with suboptimal global performance may typically exhibit superior competence within specific local subspaces.
@@ 2.3 Dynamic Ensembles
To address the rigidity of static ensembles, Dynamic Ensemble Selection (DES) has emerged [6]. Its core philosophy involves dynamically selecting the most competent subset of classifiers for each individual test sample. Existing DES methods primarily evaluate models based on a “Region of Competence” (RoC). Approaches predicated on local accuracy, such as LCA proposed by Woods et al. [7], assess models by computing classification accuracy within a local neighborhood. The KNORA series [8] adopts a more direct strategy, determining model participation based on the count of correctly classified samples in the vicinity. To circumvent the limitations of hard thresholds, RRC [25] utilizes Gaussian potential functions to assign weights based on the distance between test samples and neighbors, achieving a probabilistic competence assessment. MCB [26] introduces the concept of a “Behavioral Knowledge Space” to measure reliability by comparing the consistency of classifier output profiles in local regions. Meta-learning-based methods, such as META-DES [27], transform dynamic selection into a binary classification problem, utilizing meta-features to train a meta-classifier for competence estimation. Furthermore, FIRE-DES++ [28] incorporates online pruning strategies to enhance computational efficiency.
@@
Despite their theoretical sophistication, these DES methods encounter significant challenges regarding their dependency on external reference sets for practical deployment. Fundamentally, existing DES methods adhere to a retrieval-based paradigm: the system is required to retain the entire reference set and execute computationally expensive nearest-neighbor searches during inference. As data scales expand, time and space complexity increase linearly or even super-linearly. Moreover, in high-dimensional feature spaces, as noted by Beyer et al. [29], the discriminative power of Euclidean distance diminishes sharply (the “curse of dimensionality”), resulting in retrieved “neighbors” that are often spurious and misleading for competence assessment. In summary, while DES offers a perspective of dynamic adaptation, its structural reliance on external reference sets and neighborhood retrieval constrains its scalability to larger and higher-dimensional datasets. This limitation motivates our exploration of a novel ensemble paradigm that achieves efficient and robust dynamic integration by constructing intrinsic behavioral profiles of the models.

### v2

@@ Abstract
In the field of machine learning, ensemble learning is widely recognized as a pivotal strategy for pushing the boundaries of predictive performance. Traditional static ensemble methods typically assign weights by treating each base learner as a whole, thereby overlooking that individual models exhibit varying competence across different regions of the instance space. Dynamic Ensemble Selection (DES) was introduced to address this limitation. However, both static and dynamic approaches predominantly rely on inter-model differences as the basis for integration; this inter-model perspective neglects models’ intrinsic characteristics and often requires heavy reliance on reference sets for competence estimation. We propose the Behavioral Profiling Ensemble (BPE) framework, which introduces a model-centric integration paradigm. Unlike traditional methods, BPE constructs an intrinsic behavioral profile $\mathcal{P}_{k}$ for each model and derives aggregation weights from the deviation between a model’s response to a test instance and its established profile; in this work, we instantiate $\mathcal{P}_{k}$ with entropy-based summary statistics (e.g., mean and variance). Extensive experiments on 42 real-world datasets show that BPE-derived algorithms outperform state-of-the-art DES baselines, increasing predictive accuracy while reducing computational and storage overhead.
@@
journal: Pattern Recognition
@@ 1 Introduction
In the field of machine learning, ensemble learning improves predictive performance by combining multiple base learners and is widely used for complex classification problems. Whether in data science competitions or high-stakes industrial applications such as medical diagnosis and financial risk management [1, 2], ensemble models typically exhibit superior generalization capability and robustness compared to individual models by exploiting diversity among learners [3, 4]. For mainstream probability-based ensemble integration, strategies are generally categorized into two paradigms: global static ensembles, such as simple averaging, median averaging, or rank averaging [5, 6]; and meta-learning-based methods, such as Stacking, which learn the combination weights of base models by training a meta-classifier [7].
@@
Despite their substantial success, many global fusion rules adopt a single combination rule that is applied across the entire data distribution. However, in practical scenarios, different models often manifest distinct strengths in different regions of the instance space. A global integration perspective can therefore overlook variations in local competence. To address this issue, Dynamic Ensemble Selection (DES) and Dynamic Classifier Selection (DCS) methods have emerged [8]. Many DES/DCS methods follow a neighborhood-based validation paradigm: during the inference phase, they retrieve similar samples from a reference set and dynamically assign weights based on the models’ historical performance on these local neighbors [9, 10]. Nevertheless, such reliance on an external reference set poses challenges in deployment and may limit robustness in certain contexts. As a result, in practice (e.g., data science competitions and real-world deployments), static ensembles, especially simple probability averaging/weighting, remain more prevalent due to their simplicity and ease of deployment.
@@
To overcome these limitations, we propose a shift in perspective: moving from external reference matching to leveraging models’ intrinsic properties. This paradigm shift can be understood through an intuitive analogy. Traditional DES/DCS algorithms resemble “resume screening”: when a new task (test sample) arrives, the system reviews historical records to find individuals who performed well on similar past tasks (reference neighbors). The drawback is that for novel or outlier instances, such historical records may become less informative. In contrast, our method functions as “behavioral profiling”: rather than asking whether the individual has encountered similar tasks before, we observe inherent behavioral traits through lightweight stress testing to assign higher weights to the most reliable models.
@@
Driven by this motivation, this paper proposes the Behavioral Profiling Ensemble (BPE), a validation-free dynamic ensemble framework based on model “personality” profiling. Here, validation-free means that BPE does not require a held-out validation split to learn aggregation weights, and it avoids reference-set-based neighbor retrieval at inference time. Specifically, we avoid explicit neighborhood search on a reference set and instead introduce stochastic perturbations in the feature space to construct an instantaneous “testing field” for each model. By observing the variations in the output distributions of different base models under perturbation (e.g., entropy variation), we can characterize the models’ inherent behavioral characteristics [11, 12]. This profiling effectively differentiates between models that are robustly confident and those that exhibit high uncertainty on the current sample, thereby serving as the basis for dynamic weighting. Extensive experiments on both simulation and real-world datasets demonstrate that BPE consistently outperforms traditional methods across numerous scenarios.
The main contributions of this paper are summarized as follows:
  • 1.
    A new ensemble learning framework: This paper proposes a new ensemble learning framework, BPE, which is distinct from traditional DES/DCS. Specifically, BPE shifts the integration perspective from inter-model comparison to a model-centric one. It further introduces the notion of a model-specific behavioral profile and a deviation-based weighting principle that evaluates each model relative to itself.
  • 2.
    Key building blocks and an effective instantiation: This paper presents the basic building blocks of the BPE framework and proposes a simple instantiation, BPE-Entropy, which improves accuracy over traditional reference-set-based DES/DCS methods in many scenarios while reducing deployment-relevant storage and computational overhead.
@@ 2.1 Ensemble Learning Theory and Uncertainty Measures
The core philosophy of ensemble learning lies in compensating for the inductive bias of a single learner by combining multiple hypotheses. The “No Free Lunch” theorems proposed by Wolpert and Macready [13] state that no single learning algorithm can perform optimally across all possible data distributions, thereby motivating the use of diverse hypotheses in practice. To quantify the effectiveness of ensembles, Krogh and Vedelsby proposed the well-known “Error-Ambiguity Decomposition” theory [3]; in their formulation, the ensemble generalization error can be expressed as the average error of the base models minus an ambiguity term that captures model diversity. This theory establishes the dual objectives of ensemble learning: maximizing diversity among models while pursuing individual accuracy. From a statistical perspective, Dietterich further noted that ensemble methods can smooth decision boundaries by averaging multiple hypotheses and often reduce variance compared to individual models [4].
To accurately allocate weights in dynamic ensembles, it is essential to quantify a model’s sample-wise predictive uncertainty. Information entropy, as introduced by Shannon [14], is a fundamental measure of uncertainty for probability distributions and is widely used as a convenient proxy for confidence; however, low entropy does not necessarily imply correctness when models are miscalibrated. Grandvalet and Bengio showed in semi-supervised learning that entropy minimization can encourage confident predictions on unlabeled data [15]. In modern machine learning, perturbation- and sampling-based approaches to uncertainty estimation have become mainstream. Deep Ensembles [12] demonstrated that epistemic uncertainty can be estimated by training and combining multiple independently initialized models. Similarly, Gal and Ghahramani showed that enabling dropout at inference time provides a Bayesian approximation [11]; by aggregating multiple stochastic forward passes, one can estimate predictive uncertainty (e.g., via predictive entropy). Furthermore, Schapire’s margin theory suggests that improving classification margins—often reflected by larger gaps between top predicted probabilities—can benefit generalization [16]. Taken together, these studies motivate our BPE framework, which constructs a model-specific behavioral baseline from perturbation-induced variations in its predictive distribution and performs dynamic aggregation based on the magnitude of a model’s deviation from this baseline.
@@ 2.2 Static Ensembles
Early ensemble strategies predominantly adopted a static perspective, where the aggregation weights of base models remain fixed across the entire feature space. Depending on the generation mechanism of the base learners, these strategies are primarily categorized into homogeneous and heterogeneous ensembles; in this paper, we use “homogeneous” to refer to ensembles that generate base learners from a shared underlying model family (often decision trees) via a common mechanism (e.g., Bagging or Boosting).
In homogeneous ensembles, the Bagging algorithm proposed by Breiman [17] constructs diversity via bootstrap sampling; its prominent representative, Random Forests [18], primarily enhances performance by mitigating variance. Conversely, Boosting paradigms aim to reduce bias by iteratively correcting residuals. Notable examples include the classic AdaBoost [19] and Gradient Boosting Machines (GBM) [20], along with highly optimized implementations such as XGBoost [21], LightGBM [22], and CatBoost [23].
In heterogeneous ensembles that combine diverse algorithms (e.g., SVC [24], KNN [25], and Logistic Regression [26]), the prevailing strategy involves static fusion rules. Kittler et al. systematized fundamental combination rules, including majority voting and simple averaging [5]. To address the limited comparability of probability estimates across models (e.g., due to differing calibration), rank-based methods, such as the Borda Count applied by Ho et al. [6], mitigate these differences through rank aggregation. Furthermore, the Stacking framework [7] employs a meta-learner to learn a global (instance-invariant) combination of base models. Despite their widespread success, these static methods inherently overlook the locality of data distributions; a model with suboptimal global performance may exhibit superior competence within specific local subspaces.
@@ 2.3 Dynamic Ensembles
To address the rigidity of static ensembles, Dynamic Ensemble Selection and Dynamic Classifier Selection (DES/DCS) have been extensively studied [8]. Their core philosophy is to estimate the competence of base classifiers for each test sample and then perform dynamic selection and/or weighting accordingly. Existing DES/DCS methods primarily evaluate models based on a “Region of Competence” (RoC), which is typically constructed from a reference set. Approaches predicated on local accuracy, such as LCA proposed by Woods et al. [9], assess models by computing class-conditional accuracy within a local neighborhood. The KNORA series [10] adopts a more direct oracle-based strategy, with different variants determining model participation based on whether neighbors in the RoC are correctly classified. To circumvent the limitations of hard thresholds, RRC [27] utilizes Gaussian potential functions to assign weights based on the distance between the test sample and its reference neighbors, achieving a probabilistic competence assessment. MCB [28] introduces the concept of a behavior knowledge space to measure reliability by comparing the consistency of classifier output profiles in local regions. Meta-learning-based methods, such as META-DES [29], transform dynamic selection into a binary classification problem, utilizing meta-features to train a meta-classifier for competence estimation. Furthermore, FIRE-DES++ [30] incorporates online pruning strategies to enhance computational efficiency.
@@
Despite their theoretical sophistication, these DES/DCS methods encounter significant challenges regarding their dependency on external reference sets for practical deployment. Fundamentally, existing DES methods adhere to a retrieval-based paradigm: the system is required to retain a (potentially large) reference set and execute computationally expensive nearest-neighbor searches for each test sample during inference. As data scales expand, time and space complexity typically increase, imposing non-trivial overhead. Moreover, in high-dimensional feature spaces, as noted by Beyer et al. [31], the discriminative power of Euclidean distance can diminish (the “curse of dimensionality”), resulting in retrieved “neighbors” that may be spurious and less informative for competence assessment. In summary, while DES offers a perspective of dynamic adaptation, its structural reliance on external reference sets and neighborhood retrieval constrains its scalability to larger and higher-dimensional datasets. This limitation motivates our exploration of a novel ensemble paradigm that achieves efficient and robust dynamic integration by constructing intrinsic behavioral profiles of the models. This perspective provides a new ensemble direction: rather than relying on an external reference set or absolute statistics, we use the deviation between a model’s response and its own behavioral baseline.
3 The core intuition of BPE
In this section, we will provide some simple proofs to derive the core intuition of BPE.
Definition 1 (Models and Notations)
Let $S$ be the set of all samples in a classification task with output space $V=\{1,2,\dots,|V|\}$. We consider an ensemble of two models sharing the same input and output spaces:
  • 1.
    Let $q$ denote the Primary Model (PM).
  • 2.
    Let $q^{\prime}$ denote the Secondary Model (SM).
For any sample $s\in S$, let $q_{s}$ and $q^{\prime}_{s}$ be the output probability vectors of the Primary and Secondary models, respectively. Let $q_{s,i}$ denote the probability assigned to class $i$ by model $q$, and $k_{s}\in V$ denote the true class label of sample $s$.
Definition 2 (Partition of Sample Space)
Based on the correctness of the predictions from the Primary and Secondary models, the entire sample set $S$ is partitioned into three mutually disjoint subsets ($S=T\cup F\cup N$):
  • 1.
    $T$: Samples where the Primary Model predicts incorrectly, but the Secondary Model predicts correctly.
    $T=\{s\in S\mid\arg\max_{i\in V}(q_{s,i})\neq k_{s}\land\arg\max_{i\in V}(q^{\prime}_{s,i})=k_{s}\}$
  • 2.
    $F$: Samples where the Primary Model predicts correctly, but the Secondary Model predicts incorrectly.
    $F=\{s\in S\mid\arg\max_{i\in V}(q_{s,i})=k_{s}\land\arg\max_{i\in V}(q^{\prime}_{s,i})\neq k_{s}\}$
  • 3.
    NN: Samples where both models predict correctly, or both predict incorrectly.
    $N=\{s\in S\mid I(\arg\max q_{s}=k_{s})=I(\arg\max q^{\prime}_{s}=k_{s})\}$
Definition 3 (Theoretical Optimality of Ensemble)
Let $Q_{ens}$ be the fused predictor derived from the ensemble of $q$ and $q^{\prime}$. We define the Theoretical Optimality of the system as the state where the ensemble correctly classifies all samples that are correctly classified by at least one of the individual models, as well as the mathematically correctable subset $N_{corr}\subset N$ where both models predict incorrectly.
Formally, the ensemble achieves Theoretical Optimality if and only if:
$\forall s\in T\cup F\cup N_{corr},\quad\arg\max_{i\in V}(Q_{ens,s,i})=k_{s}$
This implies that the classification accuracy on the union set $T\cup F\cup N_{corr}$ is $100\%$.
Lemma 1 (Exchange Condition)
Let the fused model be $q_{new,s,i}=w\cdot q_{s,i}+(1-w)\cdot q^{\prime}_{s,i}$ for $w\in(0,1)$. On sample $s$, if the PM’s original prediction is $i$ (i.e., $\arg\max(q_{s})=i$), for the fused model’s prediction for class $j$ to have a higher probability than for class $i$, it must satisfy:
$w\cdot q_{s,j}+(1-w)\cdot q^{\prime}_{s,j}>w\cdot q_{s,i}+(1-w)\cdot q^{\prime}_{s,i}$
For $w\in(0,1)$, this inequality is equivalent to:
$\frac{q_{s,i}-q_{s,j}}{q^{\prime}_{s,j}-q^{\prime}_{s,i}}<\frac{1-w}{w}$
Clearly, we need SM’s prediction for $j$ to be greater than its prediction for $i$:
$q^{\prime}_{s,j}-q^{\prime}_{s,i}>0$
This is termed the exchange condition.
Furthermore, if $\arg\max(q_{new,s})=j$, then the fused model’s prediction changes from $i$ to $j$. This is termed the strict exchange condition if the exchange condition is also met.
We define the left side as the Exchange Threshold, denoted $ET(s,i\to j)$. A flip occurs only when $ET<\frac{1-w}{w}$.
This lemma implies that only when the exchange condition is met can the fused model’s prediction potentially differ from the PM’s prediction. Any case not satisfying this condition cannot change the prediction outcome. Moreover, only when the strict exchange condition is met can the change in prediction outcome be determined.
Assumption 1 (Overlap of Prediction Margins)
Typically, a model exhibits greater discrimination for samples it predicts correctly. However, to analyze the limits of static integration, we assume situations exist where this stratification is violated.
Specifically, we assume there exists at least one pair of samples, a correctable sample $t\in T$ and a vulnerable sample $f\in F$, such that the difference between the probability of the Primary Model’s incorrectly predicted class and the true class for $t$ is greater than or equal to the difference between the probability of the correctly predicted class and the second most likely class for $f$.
Formally, there exists at least one sample $t\in T$ such that for some sample $f\in F$:
$(q_{t,i_{t}}-q_{t,k_{t}})\geq(q_{f,k_{f}}-\max_{j\neq k_{f}}q_{f,j})$
where $i_{t}=\arg\max(q_{t})$ is the Primary Model’s incorrect prediction for $t$, $k_{t}$ is the true class of $t$, and $k_{f}=\arg\max(q_{f})$ is the correct prediction for $f$.
A similar condition holds for the Secondary Model $q^{\prime}$. Specifically, for the vulnerable sample $f\in F$ and the correctable sample $t\in T$, we assume:
$(q^{\prime}_{f,j_{f}}-q^{\prime}_{f,k_{f}})\geq(q^{\prime}_{t,k_{t}}-\max_{j\neq k_{t}}q^{\prime}_{t,j})$
where $j_{f}=\arg\max(q^{\prime}_{f})$ is the Secondary Model’s incorrect prediction for $f$, $k_{f}$ is the true class of $f$, and $k_{t}=\arg\max(q^{\prime}_{t})$ is the correct prediction for $t$.
Theorem 1 (Impossibility of Static Perfection)
Consider a static linear ensemble system defined by $q_{new}=w\cdot q+(1-w)\cdot q^{\prime}$ with weight $w\in(0,1)$. Under Assumption 1, where the Primary Model’s confidence in its errors is greater than or equal to its confidence in its correct predictions for certain samples (specifically, $\exists t\in T,f\in F$ such that $ET(t)\geq ET(f)$), there exists no single static weight $w$ that allows the ensemble to achieve Theoretical Optimality.
Proof 1
Theoretical Optimality requires the fused model $q_{new}$ to correctly classify all samples in $S_{opt}=T\cup F\cup N_{corr}$. Since $T\cup F\subset S_{opt}$, a strict necessary condition for Theoretical Optimality is satisfying the classification constraints for the subset $T\cup F$. Therefore, we can omit the discussion of $N_{corr}$ and focus entirely on the conflicting conditions between $T$ and $F$.
For any sample $t\in T$ (where the Primary Model is incorrect and the Secondary Model is correct), the ensemble must flip the prediction to the true class $k_{t}$. According to Lemma 1, this requires the weight ratio $\tau=\frac{1-w}{w}$ to exceed the Exchange Threshold for correction:
$\tau>\frac{q_{t,i_{t}}-q_{t,k_{t}}}{q^{\prime}_{t,k_{t}}-q^{\prime}_{t,i_{t}}}=ET(t)$ (1)
Conversely, for any sample $f\in F$ (where the Primary Model is correct and the Secondary Model is incorrect), the ensemble must preserve the true class $k_{f}$ and avoid flipping to the incorrect class $j_{f}$. According to Lemma 1, this requires $\tau$ to remain below the Exchange Threshold for error induction:
$\tau<\frac{q_{f,k_{f}}-q_{f,j_{f}}}{q^{\prime}_{f,j_{f}}-q^{\prime}_{f,k_{f}}}=ET(f)$ (2)
For a single static weight $w$ to exist, there must be a valid $\tau$ satisfying both inequalities (1) and (2) for all pairs of samples in $T$ and $F$. This implies the necessary condition:
$ET(t)<\tau<ET(f)\implies ET(t)<ET(f)$
However, Assumption 1 states that there exists at least one pair of samples $t\in T$ and $f\in F$ such that the margin stratification is violated, i.e., $ET(t)\geq ET(f)$. Substituting this into the necessary condition yields a contradiction:
$\tau>ET(t)\geq ET(f)>\tau$
This implies $\tau>\tau$, which is mathematically impossible. Consequently, no such static weight $w$ exists that can simultaneously correct $t$ and preserve $f$, rendering Theoretical Optimality unreachable.
Assumption 2 (Preservation of Margin Inversion)
Let $Q_{ens}$ be a static linear ensemble of a set of models $\{q_{i}\}_{i=1}^{M}$ with weights summing to 1. Since static weighting applies a uniform linear scaling to all output probabilities, the smoothing effect on the margins of incorrect and correct predictions is consistent.
Therefore, we assume that if the constituent models satisfy the margin violation condition (Assumption 1), the aggregated ensemble $Q_{ens}$ retains this property. Specifically, for the ensemble model, there still exists a pair of samples $t\in T,f\in F$ such that the confidence in the error remains greater than or equal to the confidence in the correct prediction:
$(Q_{ens,t,i_{t}}-Q_{ens,t,k_{t}})\geq(Q_{ens,f,k_{f}}-\max_{j\neq k_{f}}Q_{ens,f,j})$
Corollary 1 (Impossibility for Multi-Model Ensembles)
Consider an ensemble of $M$ models $\mathcal{Q}=\{q_{1},\dots,q_{M}\}$, where each model satisfies Assumption 1. Under Assumption 2, there exists no static weight distribution $\{w_{1},\dots,w_{M}\}$ that allows the ensemble to achieve Theoretical Optimality.
Proof 2
We proceed by induction on the number of models $M$.
  • 1.
    Base Case ($M=2$): This is a direct application of Theorem 1. When the margin violation condition holds ($ET(t)\geq ET(f)$), no static weight can simultaneously satisfy the conflicting requirements for correction and preservation.
  • 2.
    Inductive Step: Assume the statement holds for $k$ models. Let $Q_{k}$ be the optimal static ensemble of the first $k$ models. By the inductive hypothesis, $Q_{k}$ fails to achieve optimality. Furthermore, by Assumption 2, $Q_{k}$ retains the margin violation property.
    Now, consider adding the $(k+1)$-th model, $q_{k+1}$. The new ensemble $Q_{k+1}$ can be formulated as a pairwise integration between the existing ensemble and the new model:
    $Q_{k+1}=\alpha\cdot Q_{k}+(1-\alpha)\cdot q_{k+1}$
    Since both $Q_{k}$ (by inheritance) and $q_{k+1}$ (by definition) satisfy the margin violation condition, this mathematically reduces to the Base Case ($M=2$). Consequently, no static scalar $\alpha$ exists that can render $Q_{k+1}$ theoretically optimal without inducing errors.
By induction, Theoretical Optimality is unreachable for any finite number of models $M\geq 2$.
Theorem 2 (Algebraic Monotonicity of Ensemble Potential)
Consider an ensemble of $M$ models with weights $\mathbf{w}=[w_{1},\dots,w_{M}]$ ($w_{i}\geq 0$). For any sample $s$, let the Discriminative Margin of the $i$-th model be defined as the probability difference between the true class $k_{s}$ and the most competitive incorrect class $j$:
$\delta_{i}(s)=q_{s,k_{s}}^{(i)}-q_{s,j}^{(i)}$
The ensemble correctly classifies sample $s$ if and only if the weighted sum of margins is positive:
$\Delta_{ens}(s,\mathbf{w})=\sum_{i=1}^{M}w_{i}\delta_{i}(s)>0$
We define Model Improvement as the reduction of margin overlap (Assumption 1). Specifically:
  • 1.
    For correctable samples ($t\in T$), the incorrect prediction confidence decreases or the true class confidence increases. This means $\delta_{i}(t)$ increases (becomes less negative or turns positive).
  • 2.
    For preservation samples ($f\in F$), the correct prediction confidence increases or remains stable. This means $\delta_{i}(f)$ increases (becomes more positive).
Formally, if the system state updates such that $\delta^{\prime}_{i}(s)\geq\delta_{i}(s)$ for all $i,s$, then the global maximum accuracy is strictly non-decreasing:
$\Delta ACC^{\prime}_{max}\geq\Delta ACC_{max}$ (3)
Proof 3
Let $A(\mathbf{w})$ be the set of samples correctly classified by a specific weight vector $\mathbf{w}$.
$A(\mathbf{w})=\{s\in S\mid\sum_{i=1}^{M}w_{i}\delta_{i}(s)>0\}$
Consider the updated state where margins improve to $\delta^{\prime}_{i}(s)\geq\delta_{i}(s)$. Since all weights are non-negative ($w_{i}\geq 0$), the linearity of the ensemble summation guarantees:
$\sum_{i=1}^{M}w_{i}\delta^{\prime}_{i}(s)\geq\sum_{i=1}^{M}w_{i}\delta_{i}(s)$
This implies a strict inclusion of classification capability for any fixed $\mathbf{w}$:
  1. 1.
    If a sample was originally correct ($\sum w_{i}\delta_{i}(s)>0$), the new sum is even larger, so it remains correct.
  2. 2.
    If a sample was originally incorrect ($\sum w_{i}\delta_{i}(s)\leq 0$), the increase may push the sum above 0, correcting it.
Thus, for any candidate weight vector $\mathbf{w}$, the number of correct samples is non-decreasing: $|A^{\prime}(\mathbf{w})|\geq|A(\mathbf{w})|$. Let $\mathbf{w}^{*}$ be the optimal weight configuration for the original system. It follows that:
$\Delta ACC^{\prime}_{max}=\max_{\mathbf{w}}\frac{|A^{\prime}(\mathbf{w})|}{|S|}\geq\frac{|A^{\prime}(\mathbf{w}^{*})|}{|S|}\geq\frac{|A(\mathbf{w}^{*})|}{|S|}=\Delta ACC_{max}$
This proves that reducing the margin overlap (improving separability) strictly preserves or improves the global performance potential.
These proofs derive the core intuition behind BPE. Theorem 1 and Corollary 1 demonstrate the unavoidable theoretical defects of static integration, rendering dynamic integration necessary. However, the prevalence of static ensembles—despite their lack of theoretical advantage—raises a critical question: what form of dynamic weighting should be employed?
Theorem 2 answer this by identifying that the behavioral adjustment of a model’s internal output probabilities is the decisive factor determining the magnitude of accuracy (ACC) improvement. This insight guides a shift in the ensemble learning perspective from inter-model comparison to intra-model behavioral adjustment. Assuming that specific internal behavioral patterns form the foundation of effective ensembling (e.g., on average, a model’s output probabilities tend to be more extreme when making correct predictions), appropriately amplifying these patterns can further enhance the capability of the ensemble model.
Beyond defining the theoretical limit of ACC improvement, appropriate adjustment of internal behavior facilitates a larger segmentation space, making the search for ensemble weights significantly easier. This expanded weight determination space addresses the limitations of traditional dynamic ensemble methods (such as DES/DCS), which often rely heavily on validation sets and lack generalization, leading to distorted weight determination.
Taking the two-model case as an example, a larger difference ET(f)ET(t)ET(f)-ET(t) implies a wider feasible range for the weight ratio $\tau$, thereby simplifying the determination of valid weights.
Furthermore, the analysis of internal differences aligns with the established phenomenon that distinct models exhibit inherent behavioral disparities, often rendering their raw output probabilities directly incomparable.
Collectively, these studies provide the theoretical intuition and basis for the BPE framework to define the ensemble perspective as the behavioral gap of a model relative to its own inherent behavioral profile.